• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name LeRose

    Log entry time 20:02:28 on July 7,2010

    Entry number 330694

    keyword=Spectrometers Restored

    Spectrometers restored, Q2/3 cycled to 1000 Amps.
    Regulation resored. Septum restored.
    P0=1.131 GeV/c


    FIGURE 1